A bag contains 25 paise and 50 paise coins whose total value is Rs 30. If the number if 25 paise coins is four times that of 50 paise coins, find the number of each type of coins.
Answer by sachi(548)

the number if 25 paise coins is four times that of 50 paise coins
let 50 coin is x so 25 paisa coin is 4x
whose total value is Rs 30.
so 25*4x+x*50=30*100
or 100x+50x=3000
or 150x=3000
or x=3000/150=20
& 4x=20*4=80
the number of 25 paise coins is 80 & 50 paise coins is 20